# On balanced Hermitian threefolds with parallel Bismut torsion

### v1

@@ 5.3. The double cover and the Vaisman companions
$\pi$ induces an injective homomorphism $\pi_{\ast}:\mbox{Hol}^{\,b}(\hat{M})\rightarrow\mbox{Hol}^{\,b}(M)$. To be precise, denote by $f$ the involution on $\hat{M}$ which is in the deck transformation group of $\pi$, and fix a base point $p\in\hat{M}$ with $q=f(p)$. If $\gamma$ is a loop from $p$ to $p$, then the image $\pi(\gamma)$ is a loop from $\pi(p)$ to $\pi(p)$, and the $\nabla^{b}$-parallel transport $P_{\gamma}$ along $\gamma$ is mapped to the $\nabla^{b}$-parallel transport $P_{\pi(\gamma)}$ along $\pi(\gamma)$. Fix a path $\gamma_{0}$ from $p$ to $q$, then $x=P_{\pi(\gamma_{0})}$ is also an element in $\mbox{Hol}^{\,b}(M)$, and $\mbox{Hol}^{\,b}(M)$ is generated by the subgroup $\pi_{\ast}(\mbox{Hol}^{\,b}(\hat{M}))$ and $x$, so $\mbox{Hol}^{\,b}(M)$ contains $H$ as a subgroup of index either $1$ or $2$, depending on whether $x$ belongs to $\pi_{\ast}(\mbox{Hol}^{\,b}(\hat{M}))\cong H$ or not. Clearly, $f$ sends positive type admissible frames to negative type admissible frames, otherwise $M$ could be covered by positive type frames so $L\cong{\mathcal{O}}_{\!M}$ would be trivial, contradicting to our assumption. This means that $x$ is in the form
@@
From now on, we will assume that $(M^{3},g)$ is a compact balanced BTP threefold of middle type that is primary. The holomorphic tangent bundle $T_{\!M}$ decomposes as the orthogonal direct sum $L_{1}\oplus L_{2}\oplus L$ of complex line bundles. The decomposition does not vary holomorphically on $M$, or equivalently, each $L_{i}$ is not a holomorphic line bundle. However, if we let $F_{i}=L_{i}\oplus L$ for $i=1$, $2$, then we have
@@
[…]
@@
From this we get $\,[e_{1},e_{3}]=-\theta_{11}\!(e_{3})\,e_{1}$, $\,[e_{3},\overline{e}_{3}]=0$, $\,[e_{1},\overline{e}_{3}]=-\theta_{11}\!(\overline{e}_{3})\,e_{1}$, and
$[e_{1},\overline{e}_{1}]=\overline{\theta_{11}(\overline{e}_{1})}\,\overline{e}_{1}-\theta_{11}(\overline{e}_{1})\,e_{1}-a\,\overline{e}_{3}+a\,e_{3}.$
@@
$\alpha_{1}\ =\ \theta^{b}_{11}-\frac{a}{2}(\varphi_{3}-\overline{\varphi}_{3}),\ \ \ \ \ \ \ \ \ \alpha_{2}\ =\ \theta^{b}_{22}+\frac{a}{2}(\varphi_{3}-\overline{\varphi}_{3}).$
@@
Next we observe that the primary balanced BTP threefold $(M^{3},g)$ of middle type admits multiple complex structures compatible with the Riemannian metric $g$, or multiple orthogonal complex structures (abbreviated as OCS) in the terminology of Salamon [25]. Fix an admissible frame $e$ on $M$. Since $L$ is trivial, $e_{3}$ can be extend to a global holomorphic section $\sigma$ of $L$ so that $\nabla^{b}\sigma=0$. From now on, we will only use admissible frames on $M$ that are positive type, meaning that their $e_{3}$ coincide with $\sigma$, so $e_{3}$ is uniquely determined, while $e_{1}$, $e_{2}$ are only local and can be rotated by a function with norm $1$. Denote by $J$ the complex structure of $M$. Let us define almost complex structures $J_{1}$, $J_{2}$, $J_{3}$ on $M$ by
@@
for any positive type admissible frame $e$. Clearly, each $J_{i}$ is compatible with the metric $g$, satisfying $J\circ J_{i}=J_{i}\circ J$ and $J_{i}\circ J_{j}=J_{j}\circ J_{i}$ for any $1\leq i,j\leq 3$. There are also the almost complex structures $-J$, $-J_{i}$ compatible with $g$, which have opposite orientation with $J$. Those $J$ and $J_{i}$ generates an abelian group of $16$ elements:
@@
Case 1. $x=e_{j}$, $y=e_{k}$, $1\leq j\neq k\leq 3$. First we may ignore the case when $i\not\in\{j,k\}$, since $J_{i}$ acts the same way as $-J$ so (5.16) holds. If $i=j$, then $[J_{i}e_{i},J_{i}e_{k}]=[e_{i},e_{k}]$, so the first two terms in $N_{i}(e_{i},e_{k})$ cancel each other. Similarly, the last two terms also cancel each other, so (5.16) holds.
Case 2. $x=e_{j}$, $y=\overline{e}_{k}$, $1\leq j\neq k\leq 3$. Similarly with Case 1, we may ignore the case when $i\not\in\{j,k\}$, since $J_{i}$ acts the same way as $-J$ and we know that $N_{J}=0$. So we may assume that $i=j\neq k$ or $i=k\neq j$. In the first case, $[J_{i}e_{i},J_{i}\overline{e}_{k}]=[\sqrt{-1}e_{i},\sqrt{-1}\overline{e}_{k}]=-[e_{i},\overline{e}_{k}]$, while $[J_{i}e_{i},\overline{e}_{k}]+[e_{i},J_{i}\overline{e}_{j}]=2\sqrt{-1}[e_{i},\overline{e}_{k}]$, hence we have
@@
If $i=k\neq j$, then $[J_{i}e_{j},J_{i}\overline{e}_{i}]=-[e_{j},\overline{e}_{i}]$, while $[J_{i}e_{j},\overline{e}_{i}]+[e_{j},J_{i}\overline{e}_{i}]=-2\sqrt{-1}[e_{j},\overline{e}_{i}]$, thus
@@
We have $[e_{i},\overline{e}_{j}]=\nabla^{c}_{e_{i}}\overline{e}_{j}-\nabla^{c}_{\overline{e}_{j}}e_{i}=\sum_{k}\big{(}\overline{\theta_{jk}(\overline{e}_{i})}\,\overline{e}_{k}-\theta_{ik}(\overline{e}_{j})\,e_{k}\big{)}$, hence by (5.1) we obtain
@@
Denote by ^$\hat{\nabla}^{b}$ the Bismut connection of $(M,g,I)$. Then the above calculation shows that ^$\hat{\nabla}^{b}e_{i}=\nabla^{b}e_{i}$ for $i=1$ and $3$, while
^$\hat{\nabla}^{b}\overline{e}_{2}=\hat{\nabla}^{b}\varepsilon_{2}=\hat{\theta}^{b}_{22}\varepsilon_{2}=-\theta^{b}_{22}\overline{e}_{2}.$
Taking complex conjugation and using the fact that $\overline{\theta^{b}_{22}}=-\theta^{b}_{22}$, we get ^$\hat{\nabla}^{b}e_{2}=\nabla^{b}e_{2}$. This means that ^$\hat{\nabla}^{b}=\nabla^{b}$. Clearly, ^$\hat{\nabla}^{b}\hat{T}=0$, so $(M^{3},g,I)$ is BTP. It is not balanced, as its Gauduchon’s torsion $1$-form $\hat{\eta}=2a\psi_{3}$ which is lined up with $\varepsilon_{3}$. Since $\hat{T}^{1}_{23}=\hat{T}^{2}_{13}=0$ and $\hat{T}^{1}_{13}=\hat{T}^{2}_{23}=a$, by [35, Proposition 1.8] we know that $(M^{3},g,I)$ is a Vaisman manifold. Since $(M^{3},g,I)$ and $(M^{3},g,J)$ shares the same metric and Bismut connection, their Bismut holonomy group is the same and equals to $H=U(1)\!\times\!U(1)\!\times\!1$. Alternatively, because the Bismut connection ^$\hat{\nabla}^{b}$ preserves the orthogonal decomposition $L_{1}\oplus L_{2}\oplus L$ on $M$, the Bismut holonomy group of $(M^{3},g,I)$ is contained in $U(1)\!\times\!U(1)\!\times\!1$ thus is abelian, and it actually equals to $U(1)\!\times\!U(1)\!\times\!1$ since one computes that $\hat{R}^{b}_{1\bar{1}2\bar{2}}=\hat{R}^{b}_{2\bar{2}1\bar{1}}=-2a^{2}$ which is non-zero. ∎
@@
Since the metric of the Vaismann manifold $(M^{3},g,I)$ is not Kähler, we know that the first Betti number $b_{1}(M)>0$. By the beautiful theorem of Ornea and Verbitsky [23] on the structure of compact Vaisman manifolds, we also get: $M^{3}$ is a smooth fiber bundle over the circle $S^{1}$ with fiber being a compact Sasakian manifold $N^{5}$.
@@
Let $(M^{3},g,I)$ be a compact, non-Kähler Vaisman threefold. It is a non-balanced BTP manifold. Denote by ^$\hat{\nabla}^{b}$, $\hat{T}$, and $\hat{H}$ the Bismut connection, Chern torsion, and Bismut holonomy group of $(M^{3},g,I)$, respectively. Then by [35, Proposition 1.7], we know that there always exist admissible frames, namely a local unitary frame $\epsilon$ in $M^{3}$ such that $\nabla^{b}\epsilon_{3}=0$ with $\epsilon_{3}$ a global holomorphic vector field of constant norm, which indicates $\hat{H}\subset U(2)\times 1$, and the only non-zero torsion components are $\hat{T}^{1}_{13}=\hat{T}^{2}_{23}=a>0$. Write $\hat{H}=G\!\times\!1$ with $G\subset U(2)$. Since $\hat{H}$ is abelian, $G$ is an abelian subgroup of $U(2)$ hence is conjugate to a diagonal subgroup. This means that there exists a unitary change $\{\varepsilon_{1},\varepsilon_{2}\}$ of $\{\epsilon_{1},\epsilon_{2}\}$ so that $\hat{H}$ preserves the splitting ${\mathbb{C}}\varepsilon_{1}\oplus{\mathbb{C}}\varepsilon_{2}$. Write $\varepsilon_{i}=\sum_{j=1}^{2}U_{ij}\epsilon_{j}$, $1\leq i\leq 2$, where $U$ is a $U(2)$-valued local function, and let $\varepsilon_{3}=\epsilon_{3}$. Then under the new frame $\varepsilon$ the non-zero torsion components are still $\hat{T}^{1}_{13}=\hat{T}^{2}_{23}=a$, in other words, $\varepsilon$ is still an admissible frame of $(M^{3},g,I)$ in the sense of Definition 1.6 of [35], and the holonomy group $\hat{H}$ is contained in (hence equals to) $U(1)\!\times\!U(1)\!\times\!1$, which preserves an orthogonal decomposition of the tangent bundle into the direct sum of complex line subbundles $L_{1}\oplus L_{2}\oplus L_{3}$ where $L_{i}={\mathbb{C}}\varepsilon_{i}$.
Define an almost complex structure $J$ on $M^{3}$ by letting $J=I$ on $L_{1}\oplus L_{3}$ while letting $J=-I$ on $L_{2}$. In other words, $e$ will be a local unitary frame for $(M^{3},g,J)$ if $e_{1}=\varepsilon_{1}$, $e_{2}=\overline{\varepsilon}_{2}$ and $e_{3}=\varepsilon_{3}$. By analogous deduction as in the proofs of Lemma 5.14 and Proposition 5.15, we see that $J$ is integrable, and $(M^{3},g,J)$ is balanced BTP whose Bismut connection $\nabla^{b}$ coincides with ^$\hat{\nabla}^{b}$. Its Bismut holonomy group is equal to $U(1)\!\times\!U(1)\!\times\!1$, hence it is primary, and $(M^{3},g,I)$ is one of its Vaisman companions. This completes the proof of Proposition 5.19. ∎
@@ 5.4. Examples of Vaisman threefolds with abelian Bismut holonomy
In this subsection, let us consider some concrete examples of Vaisman threefold with abelian Bismut holonomy. The first one is a Vaisman companion of the complex nilmanifold $N^{3}$:
@@
Here $a>0$ is a constant. It is a Vaisman companion to the balanced BTP nilmanifold $N^{3}$ with structure equation $d\varphi_{1}=d\varphi_{2}=0$, $d\varphi_{3}=-a\varphi_{1\bar{1}}+a\varphi_{2\bar{2}}$.
@@
Similarly, by taking the Vaisman companions of $A_{s,t}$ or $B_{z,t}$, we get Vaisman threefolds with abelian Bismut holonomy that are deformations of the $M^{3}$ above, just as $A_{s,t}$ or $B_{z,t}$ are deformations of the balanced BTP nilmanifold $N^{3}$.

### v2

@@ 5.3. The double cover and the Vaisman companions
$\pi$ induces an injective homomorphism $\pi_{\ast}:\mbox{Hol}^{\,b}(\hat{M})\rightarrow\mbox{Hol}^{\,b}(M)$. To be precise, denote by $f$ the involution on $\hat{M}$ which is in the deck transformation group of $\pi$, and fix a base point $p\in\hat{M}$ with $q=f(p)$. If $\gamma$ is a loop from $p$ to $p$, then the image $\pi(\gamma)$ is a loop from $\pi(p)$ to $\pi(p)$, and the $\nabla^{b}$-parallel transport $P_{\gamma}$ along $\gamma$ is mapped to the $\nabla^{b}$-parallel transport $P_{\pi(\gamma)}$ along $\pi(\gamma)$. Fix a path $\gamma_{0}$ from $p$ to $q$, then $x=P_{\pi(\gamma_{0})}$ is also an element in $\mbox{Hol}^{\,b}(M)$, and $\mbox{Hol}^{\,b}(M)$ is generated by the subgroup $)\pi_{\ast}(\mbox{Hol}^{\,b}(\hat{M}$)) and $x$, so $\mbox{Hol}^{\,b}(M)$ contains $H$ as a subgroup of index either $1$ or $2$, depending on whether $x$ belongs to $\pi_{\ast}(\mbox{Hol}^{\,b}(\hat{M}))\cong H$ or not. Clearly, $f$ sends positive type admissible frames to negative type admissible frames, otherwise $M$ could be covered by positive type frames so $L\cong{\mathcal{O}}_{\!M}$ would be trivial, contradicting to our assumption. This means that $x$ is in the form
@@
From now on, we will assume that $(M^{3},g)$ is a compact balanced BTP threefold of middle type that is primary. The holomorphic tangent bundle $T_{\!M}$ decomposes as the orthogonal direct sum $L_{1}\oplus L_{2}\oplus L$ of complex line bundles. The decomposition does not vary holomorphically on $M$, or equivalently, each $L_{i}$ is not a holomorphic line bundle. However, if we let $F_{i}=L_{i}\oplus L$ for $i=1$, 22, then we have
@@
$\nabla^{c}_{\overline{e}_{j}}e_{1}=\sum_{i=1}^{3}\theta_{1i}(\overline{e}_{j})e_{i}=\theta_{11}(\overline{e}_{j})e_{1}-a\overline{\varphi}_{1}(\overline{e}_{j})e_{3}\in F_{1},\ \ \ \ \forall\ 1\leq j\leq 3.$
@@
From this we get $e_{1},e_{3}]=-\theta_{11}\!(e_{3})\,e_{1}$, $e_{3},\overline{e}_{3}]=0$, $e_{1},\overline{e}_{3}]=-\theta_{11}\!(\overline{e}_{3})\,e_{1}$, and
$e_{1},\overline{e}_{1}]=\overline{\theta_{11}(\overline{e}_{1})}\,\overline{e}_{1}-\theta_{11}(\overline{e}_{1})\,e_{1}-a\,\overline{e}_{3}+a\,e_{3}.$
@@
$\alpha_{1}\ =\ \theta^{b}_{11}-\frac{a}{2}(\varphi_{3}-\overline{\varphi}_{3}),\ \ \ \ \ \ \ \ \ \alpha_{2}\ =\ \theta^{b}_{22}+\frac{a}{2}(\varphi_{3}-\overline{\varphi}_{3}).\ \ \ \ \ \ \$
@@
Next we observe that the primary balanced BTP threefold $(M^{3},g)$ of middle type admits multiple complex structures compatible with the Riemannian metric $g$, or multiple orthogonal complex structures (abbreviated as OCS) in the terminology of Salamon [Sal95]. Fix an admissible frame $e$ on $M$. Since $L$ is trivial, $e_{3}$ can be extend to a global holomorphic section $\sigma$ of $L$ so that $\nabla^{b}\sigma=0$. From now on, we will only use admissible frames on $M$ that are positive type, meaning that their $e_{3}$ coincide with $\sigma$, so $e_{3}$ is uniquely determined, while $e_{1}$, $e_{2}$ are only local and can be rotated by a function with norm $1$. Denote by $J$ the complex structure of $M$. Let us define almost complex structures $J_{1}$, $J_{2}$, $J_{3}$ on $M$ by
@@
for any positive type admissible frame $e$. Clearly, each $J_{i}$ is compatible with the metric $g$, satisfying $J\circ J_{i}=J_{i}\circ J$ and $J_{i}\circ J_{j}=J_{j}\circ J_{i}$ for any $1\leq i,j\leq 3$. There are also the almost complex structures J-J, $J_{i}$ compatible with $g$, which have opposite orientation with $J$. Those $J$ and $J_{i}$ generates an abelian group of $16$ elements:
@@
Case 1. $x=e_{j}$, $y=e_{k}$, $1\leq j\neq k\leq 3$. First we may ignore the case when $i\not\in\{j,k\}$, since $J_{i}$ acts the same way as J-J so (5.34) holds. If $i=j$, then $J_{i}e_{i},J_{i}e_{k}]=[e_{i},e_{k}]$, so the first two terms in $N_{i}(e_{i},e_{k})$ cancel each other. Similarly, the last two terms also cancel each other, so (5.34) holds.
Case 2. $x=e_{j}$, $y=\overline{e}_{k}$, $1\leq j\neq k\leq 3$. Similarly with Case 1, we may ignore the case when $i\not\in\{j,k\}$, since $J_{i}$ acts the same way as J-J and we know that $N_{J}=0$. So we may assume that $i=j\neq k$ or $i=k\neq j$. In the first case, $J_{i}e_{i},J_{i}\overline{e}_{k}]=[\sqrt{-1}e_{i},\sqrt{-1}\overline{e}_{k}]=-[e_{i},\overline{e}_{k}]$, while $J_{i}e_{i},\overline{e}_{k}]+[e_{i},J_{i}\overline{e}_{k}]=2\sqrt{-1}[e_{i},\overline{e}_{k}]$, hence we have
@@
If $i=k\neq j$, then $J_{i}e_{j},J_{i}\overline{e}_{i}]=-[e_{j},\overline{e}_{i}]$, while $J_{i}e_{j},\overline{e}_{i}]+[e_{j},J_{i}\overline{e}_{i}]=-2\sqrt{-1}[e_{j},\overline{e}_{i}]$, thus
@@
We have $e_{i},\overline{e}_{j}]=\nabla^{c}_{e_{i}}\overline{e}_{j}-\nabla^{c}_{\overline{e}_{j}}e_{i}=\sum_{k}\big(\overline{\theta_{jk}(\overline{e}_{i})}\,\overline{e}_{k}-\theta_{ik}(\overline{e}_{j})\,e_{k}\big)$, hence by (5.1) we obtain
@@
Denote by $\hat{\nabla}^{b}$ the Bismut connection of $(M,g,I)$. Then the above calculation shows that $\hat{\nabla}^{b}e_{i}=\nabla^{b}e_{i}$ for $i=1$ and $3$, while
$\hat{\nabla}^{b}\overline{e}_{2}=\hat{\nabla}^{b}\varepsilon_{2}=\hat{\theta}^{b}_{22}\varepsilon_{2}=-\theta^{b}_{22}\overline{e}_{2}.$
Taking complex conjugation and using the fact that $\overline{\theta^{b}_{22}}=-\theta^{b}_{22}$, we get $\hat{\nabla}^{b}e_{2}=\nabla^{b}e_{2}$. This means that $\hat{\nabla}^{b}=\nabla^{b}$. Clearly, $\hat{\nabla}^{b}\hat{T}=0$, so $(M^{3},g,I)$ is BTP. It is not balanced, as its Gauduchon’s torsion $1$-form $\hat{\eta}=2a\psi_{3}$ which is lined up with $\varepsilon_{3}$. Since $\hat{T}^{1}_{23}=\hat{T}^{2}_{13}=0$ and $\hat{T}^{1}_{13}=\hat{T}^{2}_{23}=a$, by [ZZ24+, Proposition 1.8] we know that $(M^{3},g,I)$ is a Vaisman manifold. Since $(M^{3},g,I)$ and $(M^{3},g,J)$ shares the same metric and Bismut connection, their Bismut holonomy group is the same and equals to $H=U(1)\!\times\!U(1)\!\times\!1$. Alternatively, because the Bismut connection $\hat{\nabla}^{b}$ preserves the orthogonal decomposition $L_{1}\oplus L_{2}\oplus L$ on $M$, the Bismut holonomy group of $(M^{3},g,I)$ is contained in $U(1)\!\times\!U(1)\!\times\!1$ thus is abelian, and it actually equals to $U(1)\!\times\!U(1)\!\times\!1$ since one computes that $\hat{R}^{b}_{1\bar{1}2\bar{2}}=\hat{R}^{b}_{2\bar{2}1\bar{1}}=-2a^{2}$ which is non-zero. ∎
@@
Since the metric of the Vaismann manifold $(M^{3},g,I)$ is not Kähler, we know that the first Betti number $b_{1}(M)>0$. By the beautiful theorem of Ornea and Verbitsky [OV03] on the structure of compact Vaisman manifolds, we also get: $M^{3}$ is a smooth fiber bundle over the circle $S^{1}$ with fiber being a compact Sasakian manifold $N^{5}$.
@@
Let $(M^{3},g,I)$ be a compact, non-Kähler Vaisman threefold. It is a non-balanced BTP manifold. Denote by $\hat{\nabla}^{b}$, $\hat{T}$, and $\hat{H}$ the Bismut connection, Chern torsion, and Bismut holonomy group of $(M^{3},g,I)$, respectively. Then by [ZZ24+, Proposition 1.7], we know that there always exist admissible frames, namely a local unitary frame $\epsilon$ in $M^{3}$ such that $\nabla^{b}\epsilon_{3}=0$ with $\epsilon_{3}$ a global holomorphic vector field of constant norm, which indicates $\hat{H}\subset U(2)\times 1$, and the only non-zero torsion components are $\hat{T}^{1}_{13}=\hat{T}^{2}_{23}=a>0$. Write $\hat{H}=G\!\times\!1$ with $G\subset U(2)$. Since $\hat{H}$ is abelian, $G$ is an abelian subgroup of $U(2)$ hence is conjugate to a diagonal subgroup. This means that there exists a unitary change $\{\varepsilon_{1},\varepsilon_{2}\}$ of $\{\epsilon_{1},\epsilon_{2}\}$ so that $\hat{H}$ preserves the splitting ${\mathbb{C}}\varepsilon_{1}\oplus{\mathbb{C}}\varepsilon_{2}$. Write $\varepsilon_{i}=\sum_{j=1}^{2}U_{ij}\epsilon_{j}$, $1\leq i\leq 2$, where $U$ is a $U(2)$-valued local function, and let $\varepsilon_{3}=\epsilon_{3}$. Then under the new frame $\varepsilon$ the non-zero torsion components are still $\hat{T}^{1}_{13}=\hat{T}^{2}_{23}=a$, in other words, $\varepsilon$ is still an admissible frame of $(M^{3},g,I)$ in the sense of Definition 1.6 of [ZZ24+], and the holonomy group $\hat{H}$ is contained in (hence equals to) $U(1)\!\times\!U(1)\!\times\!1$, which preserves an orthogonal decomposition of the tangent bundle into the direct sum of complex line subbundles $L_{1}\oplus L_{2}\oplus L_{3}$ where $L_{i}={\mathbb{C}}\varepsilon_{i}$.
Define an almost complex structure $J$ on $M^{3}$ by letting $J=I$ on $L_{1}\oplus L_{3}$ while letting J=IJ=-I on $L_{2}$. In other words, $e$ will be a local unitary frame for $(M^{3},g,J)$ if $e_{1}=\varepsilon_{1}$, $e_{2}=\overline{\varepsilon}_{2}$ and $e_{3}=\varepsilon_{3}$. By analogous deduction as in the proofs of Lemma 5.17 and Proposition 5.18, we see that $J$ is integrable, and $(M^{3},g,J)$ is balanced BTP whose Bismut connection $\nabla^{b}$ coincides with $\hat{\nabla}^{b}$. Its Bismut holonomy group is equal to $U(1)\!\times\!U(1)\!\times\!1$, hence it is primary, and $(M^{3},g,I)$ is one of its Vaisman companions. This completes the proof of Proposition 5.22. ∎
@@ 5.4. Examples of Vaisman threefolds with abelian Bismut holonomy
In this subsection, let us consider some concrete examples of Vaisman threefold with abelian Bismut holonomy. The first one is a Vaisman companion of the complex nilmanifold determined by $N^{3}$, since the corresponding simply-connected Lie group $G$ admits uniform lattice as shown in Subsection 5.2:
@@
Here $a>0$ is a constant. It is a Vaisman companion to the balanced BTP nilmanifold determined by $N^{3}$ with the structure equation $d\varphi_{1}=d\varphi_{2}=0$, $d\varphi_{3}=-a\varphi_{1\bar{1}}+a\varphi_{2\bar{2}}$.
@@
Similarly, by taking Vaisman companion of the solvmanifold determined by $A_{s,t}$ or $B_{z,t}$, when the corresponding simply-connected Lie groups have uniform lattices (when the parameter is rational), we get examples of compact Vaisman threefolds with abelian Bismut holonomy. See [AO20] for a more detailed discussion on such Vaisman manifolds.